OpenStudy (anonymous):

A curve is such that \[{dy \over dx} = {3 \over (1+2x)^2}\] and the point (1, 1/2) lies on the curve. Find the equation of the curve.

OpenStudy (anonymous):

Isn't it you're meant to integrate first?

OpenStudy (lalaly):

u integrate it to find equation of thee curve,, Then u substitute the points to find the constant of integration

OpenStudy (lalaly):

do u know how to integrate this?

OpenStudy (anonymous):

\[dy \div 3 = dx \div (1+2x)^{2}\] Now integrate both sides. Put t = (1+2x). Now the integration should be easy. Find the constant of integration by putting (1,0.5) in the curve. This should give you the answer.

OpenStudy (anonymous):

Hmm never heard of this way... How do you do it?

OpenStudy (lalaly):

\[\huge{y=\int\limits{\frac{3}{(1+2x)^3}dx}}\]

OpenStudy (anonymous):

If you put t = 1+2x, then dt/2 = dx (by differentiating both sides). Now, the RHS becomes: 1/2 dt / (t^2). Now integrate it. It's called the variable separation method - Separate the variables involved, one variable on either side of the equation and then integrate both sides.

OpenStudy (lalaly):

use substitution let u=1+2x du=2dx so the integral becomes \[\int\limits{\frac{3}{u^3} \times \frac{du}{2}}\]

OpenStudy (anonymous):

Hmm Still lost :/

OpenStudy (lalaly):

simple integration u just move the dx to the other side so u can integrate with respect to x

OpenStudy (anonymous):

I got \[{ 3(1+2x)^{-1} \over -1} +c\]

OpenStudy (anonymous):

But it says that is the first step... and that you have to divide by 2 to equate it to y. How come?

OpenStudy (anonymous):

@lalaly

OpenStudy (lalaly):

yeah do u know integration by substitution?

OpenStudy (anonymous):

No..

OpenStudy (lalaly):

ok let u=1+2x so du=2dx so dx=du/2 so u substitute thos values to make it easier to integrate\[\huge{\int\limits{\frac{3}{u^3} \times \frac{du}{2}}}\]

OpenStudy (lalaly):

now u integrate that with respect to u then resubstitute the u=1+2x when ur done

OpenStudy (anonymous):

How?

OpenStudy (lalaly):

i suggest u learn the basic methods of integration, by substitution, by parts .... etc... i cant do this for u if u dont know it at all

OpenStudy (anonymous):

I learned it a different way...

OpenStudy (anonymous):

@lalaly

OpenStudy (lalaly):

how?

OpenStudy (anonymous):

\[\int\limits 3 (1 +2x)^{-2}\] \[3 \int\limits (1+2x)^{-2}\] \[3 \int\limits (1+2x)^{-2} \times 1/2\] \[3/2 (1+2x)^{-1}\]

OpenStudy (lalaly):

okthats right, then nevermind :) but remember for an indefinite integral u add a constant so\[\huge{y=\frac{3}{2}\frac{1}{1+2x)^2}+C}\]

OpenStudy (anonymous):

Oh, yeh. But how do I substitute the values?

OpenStudy (lalaly):

to the power of -2 not -1 .. in ur last post

OpenStudy (anonymous):

But it's integrated.. and it says that's right..?

OpenStudy (lalaly):

for the point(1,1/2) \[\huge{\frac{1}{2}=\frac{3}{2(1+2(1))^2}}+C\] now just solve for C

OpenStudy (anonymous):

Ok, thanks :)

OpenStudy (anonymous):

But it is ^-1

OpenStudy (lalaly):

no its -2 .. because it was -3 integrating that u just add 1 to the power which becomes -2

OpenStudy (anonymous):

But it was originally ^-2

OpenStudy (lalaly):

oh latex makes it look like a 3,, im sorry i thought it was 3 never mind lol

OpenStudy (anonymous):

That's Ok. Thanks!